OpenStudy (briellshaw):

Joshua has 16 bad girls as his friends he kissed 1/4 of them how many left does he have to kiss? 12 13 1 4

Mehek (mehek14):

divide 16/4

OpenStudy (briellshaw):

12 i know

Mehek (mehek14):

ok lol

Nnesha (nnesha):

is that your original question ??

Mehek (mehek14):

no that's the first step of solving it

Mehek (mehek14):

she went directly to the answer

OpenStudy (briellshaw):

1/4=4 he there are 16 of them for him to kiss

Mehek (mehek14):

16/4 = 4 1/4 = 4 16 - 4 = 12

OpenStudy (briellshaw):

16-4=12 duh

Nnesha (nnesha):

there is no need to change the wording of question that is really inappropriate

OpenStudy (briellshaw):

i just blocked you bye bye

OpenStudy (briellshaw):

imma unblock you sorry that was my lil bro talking

Nnesha (nnesha):

thanks for doing that and yes BYE

Nnesha (nnesha):

thanks for blocking* me**

OpenStudy (briellshaw):

im sorry hes really rude i didnt know he was on my account

OpenStudy (briellshaw):

im sorry i just did it

Mehek (mehek14):

wait did your brother post the question?

OpenStudy (briellshaw):

yeah

OpenStudy (briellshaw):

he did it while i was in the bathroom

OpenStudy (briellshaw):

@Nnesha im sorry for my lil bro's action's

Nnesha (nnesha):

it's okay take care of ur acc